OpenStudy (anonymous):

r=2/(5l)^-1/2 for L.

myininaya (myininaya):

I don't see L to solve for. Does l=L?

myininaya (myininaya):

\[r=2/(\sqrt{5L})\]

myininaya (myininaya):

is this right?

myininaya (myininaya):

\[r=2/(5L)^{-1/2}\]

myininaya (myininaya):

which one is right?

OpenStudy (anonymous):

\[ r= 2/(5L)^{-1/2}\]

myininaya (myininaya):

ok

myininaya (myininaya):

make that exponent positive by putting whatever the base of that exponent is in the numerator therefore making the exponent positive

myininaya (myininaya):

like r=2(5L)^{1/2}

myininaya (myininaya):

now divide by 2 giving u r/2=(5L)^(1/2)

myininaya (myininaya):

to get rid of the square root square both sides giving you (r/2)^2=5L

myininaya (myininaya):

Divide both sides by 5 giving you (1/5)*(r/2)^2=L

OpenStudy (anonymous):

okay thanks

myininaya (myininaya):

your welcome